# Memory effects in a dynamical decoupling process by parity kicks

### v1

@@ IV.3 Effect and performance of control
$\mathrm{or}$ $P[\rho_{S}(0)]=1-D[\rho_{S}(t_{f}),\rho_{\mathrm{tar}}(t_{f})]$ (52)
brought by the control where $\rho_{T(t_{f},0)}$ [$\rho_{S}(t_{f})$] represents the final dynamical map (final state) without or with control. For example, for perfect decoupling, $\Delta P=\tilde{P}-\bar{P}=E$ where $\tilde{P}=1-D[\rho_{\tilde{T}(t_{f},0)},\rho_{T_{\mathrm{tar}}(t_{f},0)}]=1$ is the performance with control, and $\bar{P}=1-D[\rho_{\bar{T}(t_{f},0)},\rho_{T_{\mathrm{tar}}(t_{f},0)}]$ is that without control. The interpretation approximately holds if $\tilde{T}(t_{f},0)\approx T_{\mathrm{tar}}(t_{f},0)$ and $D[\rho_{\tilde{T}(t_{f},0)},\rho_{\bar{T}(t_{f},0)}]\gg D[\rho_{\tilde{T}(t_{f},0)},\rho_{T_{\mathrm{tar}}(t_{f},0)}]$. In terms of quantum states, these conditions are $\tilde{\rho}_{S}(t_{f})\approx\rho_{\mathrm{tar}}(t_{f})$ and $D[\tilde{\rho}_{S}(t_{f}),\bar{\rho}_{S}(t_{f})]\gg D[\tilde{\rho}_{S}(t_{f}),\rho_{\mathrm{tar}}(t_{f})]$. In the case of ineffective dynamical decoupling, i.e., $\tilde{T}(t_{f},0)\neq T_{\mathrm{tar}}(t_{f},0)$ or $\tilde{\rho}_{S}(t_{f})\neq\rho_{\mathrm{tar}}(t_{f})$ the above interpretation may still hold if the control makes the uncontrolled final dynamical map (state) move toward the target one.
@@ V Example
V.1 Model and control
@@
In this section, we illustrate our theory with a quantum Rabi model where a qubit interacts with a dissipative harmonic oscillator in a Markovian reservoir at temperature $T$. The parity kicks are implemented on the qubit, i.e., the system (denoted by $S$), to protect it from the influence of the environment (the dissipative oscillator, denoted by $E$). The Hamiltonian of the qubit and oscillator is given by
@@
The parity kicks are chosen as $K_{P}=\sigma_{z}$, which is assumed to be ideal. It leads to an instantaneous phase change in the qubit’s quantum state. Staring from $t=0$, the controlled dynamics can be described by
@@
where $H_{1}(t)=\sum_{j=1}^{n}\delta(t-t_{j})H_{P}$ that realizes $K_{P}=\sigma_{z}$ at $t_{j}=j\tau$. It is easy to seen that conditions Eq. (5) and (6) are satisfied for our scheme. Therefore, the qubit can be dynamically decoupled from its environment after one or more cycles of parity kicks if $t_{c}$ is short enough. Staring from the condition $\rho_{SE}(0)=\rho_{S}(0)\otimes\rho_{E}^{\mathrm{th}}$ with arbitrary $\rho_{S}(0)$, the Choi-Jamiółkowski matrix of the target final dynamical map $T_{\mathrm{tar}}(t_{f},0)$ is given by
@@
V.2 Commutation relation between $\mathcal{K}_{P}$ and $\bar{T}(t_{b},t_{a})$
To investigate the role of memory effects with our theory, the condition $\mathcal{K}_{P}\bar{T}(t_{b},t_{a})=\bar{T}(t_{b},t_{a})\mathcal{K}_{P}$ is required. In general, the dynamical map $\bar{T}(t_{b},t_{a})$ depends on the master equation (54) and the environment initial state $\rho_{E}(t_{I})$. With a thermal environment initial state Eq. (56) and $K_{P}=\sigma_{z}$, we find that $\mathcal{K}_{P}$ indeed commutes with $\bar{T}(t_{b},t_{a})$. This condition can be straightforwardly verified by numerically simulating Eq. (54) before and after the kick (with different qubit initial states), or directly calculating and comparing the Choi-Jamiółkowski matrices $\rho_{\mathcal{K}_{P}\bar{T}(t_{b},t_{a})}$ and $\rho_{\bar{T}(t_{b},t_{a})\mathcal{K}_{P}}$ by simulations. Besides numerical evidences, we provide an analytical explanation based on the forms of Eq. (54) and Eq. (56) and the properties Eq. (5) and (6). In our work, the evolution of the qubit in the absence of control could be understood as $\rho_{S}(t_{b})=\bar{T}(t_{b},t_{a})\rho_{S}(t_{a})=\text{Tr}_{E,R}[e^{-iH_{S\!E\!R}(t_{b}-t_{a})}\rho_{S}(t_{a})\otimes\rho_{E}^{\mathrm{th}}\otimes\rho_{R}^{\mathrm{th}}e^{iH_{S\!E\!R}(t_{b}-t_{a})}]$ where
@@ V.2 Commutation relation between $\mathcal{K}_{P}$ and $\bar{T}(t_{b},t_{a})$
represents the Hamiltonian of the qubit, the oscillator, a Markovian reservoir (denoted by R) and their interactions. The reservoir is in a thermal state at temperature $T$, which is given by
$\rho_{R}^{\mathrm{th}}=\prod_{j}\sum_{n_{j}=0}^{\infty}\frac{\bar{n}(\omega_{j})}{[1+\bar{n}(\omega_{j})]^{n_{j}+1}}|n_{j}\rangle\langle n_{j}|.$ (61)
Then the $E\!-\!R$ initial state can be concisely expressed as
$\rho_{E}^{\mathrm{th}}\otimes\rho_{R}^{\mathrm{th}}$ $=$ $\sum p_{\mathbf{n}}|\mathbf{n}\rangle\langle\mathbf{n}|$ (62)
where $|\mathbf{n}\rangle=|n_{0},n_{1},n_{2}\cdots\rangle$, $p_{\mathbf{n}}=p_{n_{0},n_{1},n_{2}\cdots}$ is the diagonal element in the basis $|\mathbf{n}\rangle$, and all the non-diagonal elements are zero. Now we divide the $S\!-\!E\!-\!R$ Hamiltonian into $H_{+}=H_{S}+H_{E}+H_{R}+H_{E\!R}$ and $H_{-}=H_{S\!E}$ such that $\sigma_{z}H_{+}\sigma_{z}=H_{+}$ and $\sigma_{z}H_{-}\sigma_{z}=-H_{-}$. Moreover, it is seen that the action $H_{+}|\tilde{n}\rangle$ ($\langle\tilde{n}|H_{+}$) turns $|\mathbf{n}\rangle$ ($\langle\mathbf{n}|$) into bases with the same total excitation number (or annihilate it), e.g., the term $g_{1}ab_{1}^{{\dagger}}$ in $H_{+}$ turns $|n,n_{1},n_{2}\cdots\rangle$ into $|n_{0}-1,n_{1}+1,n_{2}\cdots\rangle$ (up to a coefficient) if $n_{0}\geqslant 1$. In contrast, $H_{-}|\mathbf{n}\rangle$ ($\langle\mathbf{n}|H_{-}$) turns $|\mathbf{n}\rangle$ ($\langle\mathbf{n}|$) into bases with one more or one less excitations (or annihilate it). Now the evolution of the qubit in the absence of control can be described by
$\rho_{S}(t_{b})$ $=$ $\bar{T}(t_{b},t_{a})\rho_{S}(t_{a})$ (63)
$=$ $\sum_{\tilde{n}^{\prime}}\langle\tilde{n}^{\prime}|U_{s}\rho_{S}(t_{a})\otimes(\sum_{\bar{n}}p_{\tilde{n}}|\mathbf{n}\rangle\langle\mathbf{n}|)U_{s}^{{\dagger}}|\tilde{n}^{\prime}\rangle$
where $U_{s}=I-i(H_{+}+H_{-})(t_{b}-t_{a})-(H_{+}+H_{-})^{2}(t_{b}-t_{a})^{2}/2+\cdots$ is the Taylor series of $e^{-iH_{S\!E\!R}(t_{b}-t_{a})}$. Note that Eq. (63) can be expressed as a summation of infinite terms, i.e., $\rho_{S}(t_{b})=\sum_{l}\rho_{S}^{(l)}$. Any term $\rho_{S}^{(l)}$ satisfies
$\rho_{S}^{(l)}\propto\langle\mathbf{n}^{\prime}|H_{\alpha}\!\cdots\!H_{\alpha}\rho_{S}(t_{a})\otimes|\mathbf{n}\rangle\langle\mathbf{n}|H_{\alpha}\!\cdots\!H_{\alpha}|\mathbf{n}^{\prime}\rangle$ (64)
where $\alpha$ represents $+$ or -- and the number of $H_{\alpha}$ in Eq. (64) ranges from 0 to $\infty$. The coefficient in Eq. (64) is ignored for simplicity now and later.
It can be verified that if the number of $H_{-}$ in Eq. (64) is odd, then $\rho_{S}^{(l)}=0$. For example, there are three $H_{-}$ in the term $H_{-}H_{+}\rho_{S}(t_{a})\otimes|\mathbf{n}\rangle\langle\mathbf{n}|H_{-}H_{-}H_{+}$. See that $H_{-}H_{+}$ turns $|\mathbf{n}\rangle$ into bases with 1 more or 1 less excitations (denoted by $|\mathbf{n},\pm 1\rangle$, or annihilate it), whereas $H_{-}H_{-}H_{+}$ turns $\langle\mathbf{n}|$ into bases with 2 more, 2 less or invariant excitations (denoted by $\langle\mathbf{n},+0,\pm 2|$, or annihilated it). Then $\langle\mathbf{n}^{\prime}|\hat{\rho}_{S}\otimes|\mathbf{n},\pm 1\rangle\langle\mathbf{n},+0,\pm 2|\mathbf{n}^{\prime}\rangle=0$ where $\hat{\rho}_{S}$ is transformed from $\rho_{S}(t_{a})$ by the Hamiltonians on the left-hand side of $\rho_{S}(t_{a})$. Similarly, any term with an odd number of $H_{-}$ turns $|\mathbf{n}\rangle$ and $\langle\mathbf{n}|$ into bases with different excitation numbers or annihilate one (or two) of them. Then, any term obtained by expanding Eq. (64), if not annihilated, satisfies
$\langle\mathbf{n}^{\prime}|\bar{\rho}_{S}(t_{a})\otimes|\mathbf{n},+n_{\mathrm{odd}}\rangle\langle\mathbf{n},+n_{\mathrm{even}}|\mathbf{n}^{\prime}\rangle$ $=$ $0,$ (65)
$\mathrm{or}\ \langle\mathbf{n}^{\prime}|\bar{\rho}_{S}(t_{a})\otimes|\mathbf{n},+n_{\mathrm{even}}\rangle\langle\mathbf{n},+n_{\mathrm{odd}}|\mathbf{n}^{\prime}\rangle$ $=$ 00 0 (66)
where $n_{\mathrm{odd}}=\pm 1,\pm 3,\pm 5\cdots$ and $n_{\mathrm{even}}=0,\pm 2,\pm 4\cdots$. So $\rho_{S}^{(l)}=0$ is guaranteed. On the other hand, if the number of $H_{-}$ is even in Eq. (64), there is
$\sigma_{z}H_{\alpha}\!\cdots\!H_{\alpha}\rho_{S}(t_{a})\otimes|\mathbf{n}\rangle\langle\mathbf{n}|H_{\alpha}\!\cdots\!H_{\alpha}\sigma_{z}$ (67)
$=$ $(-1)^{n_{\mathrm{even}}}H_{\alpha}\!\cdots\!H_{\alpha}\sigma_{z}\rho_{S}(t_{a})\otimes|\mathbf{n}\rangle\langle\mathbf{n}|\sigma_{z}H_{\alpha}\!\cdots\!H_{\alpha}$
$=$ $H_{\alpha}\!\cdots\!H_{\alpha}\sigma_{z}\rho_{S}(t_{a})\sigma_{z}\otimes|\mathbf{n}\rangle\langle\mathbf{n}|H_{\alpha}\!\cdots\!H_{\alpha}$
since $\sigma_{z}H_{-}=-H_{-}\sigma_{z}$ and $\sigma_{z}H_{+}=H_{+}\sigma_{z}$. Now we know that every term in Eq. (63) with an even number of $H_{-}$ satisfies Eq. (67) for any system initial state. By inserting Eq. (67) back into Eq. (63), $\mathcal{K}_{P}\bar{T}(t_{b},t_{a})=\bar{T}(t_{b},t_{a})\mathcal{K}_{P}$ is obtained.
@@
Following similar analysis or through numerical simulations, it is able to verify that $\mathcal{K}_{P}\bar{T}(t_{b},t_{a})=\bar{T}(t_{b},t_{a})\mathcal{K}_{P}$ might also holds for more related models. For example, a dissipative (or undamped) JC model, our model where the oscillator’s initial state satisfies $\rho_{E}(t_{I})=\sum_{n_{0}=0}^{\infty}p_{n_{0}}|n_{0}\rangle\langle n_{0}|$ (not limited to thermal states), a qubit coupled to multiple dissipative (or undamped) oscillators, a qubit coupled to a set of interacting oscillators, etc. Moreover, as known from the pseudomode theory [45, 46] and its recent progresses [49, 48, 47], the reduced dynamics of an open quantum system with a complex environment (infinite modes) is equivalent to that with a simpler environment (finite modes each subject to a Lindblad evolution) under certain conditions. Therefore, it is possible to apply our theory to more complex environments [50] to study the role of memory effects in a parity kick process.

### v2

@@ IV.3 Effect and performance of control
$\mathrm{or}\ \$$P[\rho_{S}(0)]=1-D[\rho_{S}(t_{f}),\rho_{\mathrm{tar}}(t_{f})]$ (50)
brought by the control where $\rho_{T(t_{f},0)}$ $\rho_{S}(t_{f})$] represents the final dynamical map (final state) without or with control. For example, for perfect decoupling, $\Delta P=\tilde{P}-\bar{P}=E$ where $\tilde{P}=1-D[\rho_{\tilde{T}(t_{f},0)},\rho_{T_{\mathrm{tar}}(t_{f},0)}]=1$ is the performance with control, and $\bar{P}=1-D[\rho_{\bar{T}(t_{f},0)},\rho_{T_{\mathrm{tar}}(t_{f},0)}]$ is that without control. The interpretation approximately holds if $\tilde{T}(t_{f},0)\approx T_{\mathrm{tar}}(t_{f},0)$ and $D[\rho_{\tilde{T}(t_{f},0)},\rho_{\bar{T}(t_{f},0)}]\gg D[\rho_{\tilde{T}(t_{f},0)},\rho_{T_{\mathrm{tar}}(t_{f},0)}]$. In terms of quantum states, these conditions are $\tilde{\rho}_{S}(t_{f})\approx\rho_{\mathrm{tar}}(t_{f})$ and $D[\tilde{\rho}_{S}(t_{f}),\bar{\rho}_{S}(t_{f})]\gg D[\tilde{\rho}_{S}(t_{f}),\rho_{\mathrm{tar}}(t_{f})]$. In the case of ineffective dynamical decoupling, i.e., $\tilde{T}(t_{f},0)\neq T_{\mathrm{tar}}(t_{f},0)$ or $\tilde{\rho}_{S}(t_{f})\neq\rho_{\mathrm{tar}}(t_{f})$ the above interpretation may still hold if the control makes the uncontrolled final dynamical map (state) move toward the target one.
@@ V Example
In this section, we propose a class of system-environment Hamiltonians with certain environment initial states that support the commutation relation between a parity kick and the dynamical map without control. After that, we illustrate our theory with a specific example in which the qubit in a dissipative quantum Rabi model is controlled by parity kicks in different regimes.
V.1 Commutation relation between $\mathcal{K}_{P}$ and $\bar{T}(t_{b},t_{a})$ for a class of models
To investigate the role of memory effects in parity kicks with our theory, $\mathcal{K}_{P}\bar{T}(t_{b},t_{a})=\bar{T}(t_{b},t_{a})\mathcal{K}_{P}$ is required. The existence of this relation depends on the properties of the parity kick given by Eq. (5) and (6), and that of the dynamical map $\bar{T}(t_{b},t_{a})$ which is a function of $H_{0}=H_{S}+H_{E}+H_{S\!E}$ and the environment initial state $\rho_{E}(t_{a})=\rho_{E}(t_{I})$ defined in Eq. (17). Assume that the environment consists of $m$ subsystems ($m\geqslant 1$) whose initial state has the form $\rho_{E}(t_{a})=\sum\rho_{\mathbf{n}\mathbf{n^{\prime}}}|\mathbf{n}\rangle\langle\mathbf{n^{\prime}}|$ where $|\mathbf{n}\rangle=|n_{1},n_{2}\cdots n_{m}\rangle$ ($|\mathbf{n}^{\prime}\rangle=|n_{1}^{\prime},n_{2}^{\prime}\cdots n_{m}^{\prime}\rangle$) and $n_{j=1,2,\cdots m}\geqslant 0$ is the excitation number of the $j$th subsystem. When $|\mathbf{n}\rangle=|\mathbf{n}^{\prime}\rangle$, $\rho_{\mathbf{n}\mathbf{n}^{\prime}}$ represents the diagonal element, otherwise, when $\langle{\mathbf{n}}|\mathbf{n}^{\prime}\rangle=0$, $\rho_{\mathbf{n}\mathbf{n^{\prime}}}$ represents the off-diagonal element. We consider a class of models with the following properties.
(1) At $t_{a}=0$, for every nonzero matrix element $p_{\mathbf{n}\mathbf{n}’}$ in $\rho_{E}(t_{a})$, the total excitation numbers in $|\mathbf{n}\rangle$ and $\langle\mathbf{n}^{\prime}|$ always have the same parity, i.e., $n_{\mathrm{ket}}=\sum_{j=1}^{m}{n_{j}}$ and $n_{\mathrm{bra}}=\sum_{j=1}^{m}{n_{j}^{\prime}}$ are both even or odd. A typical example is a fully decohered state $\sum p_{\mathbf{n}\mathbf{n}}|\mathbf{n}\rangle\langle\mathbf{n}|$ where $n_{\mathrm{ket}}=n_{\mathrm{bra}}$.
(2) The action of the environment Hamiltonian $H_{E}$ on $|\mathbf{n}\rangle$ turns $|\mathbf{n}\rangle$ into bases whose total excitation numbers have the same parity with $n_{\textrm{ket}}$ (or annihilate $|\mathbf{n}\rangle$), that is, $H_{E}$ preserves the parity of the total excitation number in $|\mathbf{n}\rangle$. Consequently, the environment initial state after $t_{a}=0$, i.e., $\rho_{E}(t_{a})=e^{-iH_{E}t_{a}}\rho_{E}(0)e^{iH_{E}t_{a}}$, also satisfies the first property. A typical example is a Hamiltonian $H_{E}$ that conserves the total excitation number.
(3) The action of the interaction Hamiltonian $H_{S\!E}$ on $|\mathbf{n}\rangle$ turns $|\mathbf{n}\rangle$ into bases whose total excitation numbers have the opposite parity with $n_{\textrm{ket}}$ (or annihilate $|\mathbf{n}\rangle$), that is, $H_{S\!E}$ changes the parity of the total excitation number in $|\mathbf{n}\rangle$. For example, a Hamiltonian $H_{S\!E}$ that increases or decreases the total excitation number in $|\mathbf{n}\rangle$ by 1.
In the following, we show that $\mathcal{K}_{P}$ commutes with the dynamical map $\bar{T}(t_{b},t_{a})$ with the above properties. We first divide the Hamiltonian $H_{0}$ into $H_{+}=H_{S}+H_{E}$ and $H_{-}=H_{S\!E}$ such that
$K_{P}H_{+}$ $=$ $H_{+}K_{P}$ (51)
$K_{P}H_{-}$ $=$ $-H_{-}K_{P}$ (52)
due to Eq. (5) and (6). Meanwhile, $H_{+}$ ($H_{-}$) preserves (changes) the parity of the total excitation number in $|\mathbf{n}\rangle$. Then the evolution of the system in the absence of control can be described by
$\rho_{S}(t_{b})$ $=$ $\bar{T}(t_{b},t_{a})\rho_{S}(t_{a})$ (53)
$=$ $\sum_{\mathbf{n}^{\prime\prime}}\langle\mathbf{n}^{\prime\prime}|U_{0}\rho_{S}(t_{a})\otimes(\sum\rho_{\mathbf{n}\mathbf{n^{\prime}}}|\mathbf{n}\rangle\langle\mathbf{n^{\prime}}|)U_{0}^{{\dagger}}|\mathbf{n}^{\prime\prime}\rangle\quad\quad$
where $U_{0}=I-i(H_{+}+H_{-})(t_{b}-t_{a})-(H_{+}+H_{-})^{2}(t_{b}-t_{a})^{2}/2+\cdots$ is the Taylor series of $e^{-iH_{0}(t_{b}-t_{a})}$ and $|\mathbf{n}^{\prime\prime}\rangle=|n_{1}^{\prime\prime},n_{2}^{\prime\prime}\cdots n_{m}^{\prime\prime}\rangle$. Moreover, $n_{\mathrm{ket}}$ and $n_{\mathrm{bra}}$ have the same parity. Note that Eq. (53) can be expressed as a summation of infinite terms, i.e., $\rho_{S}(t_{b})=\sum_{l}\rho_{S}^{(l)}$. Any term $\rho_{S}^{(l)}$ satisfies
$\rho_{S}^{(l)}\propto\langle\mathbf{n}^{\prime\prime}|H_{\alpha}\!\cdots\!H_{\alpha}\rho_{S}(t_{a})\otimes|\mathbf{n}\rangle\langle\mathbf{n}^{\prime}|H_{\alpha}\!\cdots\!H_{\alpha}|\mathbf{n}^{\prime\prime}\rangle\quad\quad$ (54)
where $\alpha$ represents ++ or - and the number of $H_{\alpha}$ in Eq. (54) ranges from 0 to $\infty$. The coefficient in Eq. (54) is ignored for simplicity now and later.
It can be verified that if the number of $H_{-}$ in Eq. (54) is odd, then $\rho_{S}^{(l)}=0$. For example, there are three $H_{-}$ in the term $H_{-}H_{+}\rho_{S}(t_{a})\otimes|\mathbf{n}\rangle\langle\mathbf{n}^{\prime}|H_{-}H_{-}H_{+}$, See that the action of $H_{-}H_{+}$ on $|\mathbf{n}\rangle$ increases the number of excitations by $n_{\mathrm{odd}}=\pm 1,\pm 3,\pm 5\cdots$, or annihilate it (bases after action denoted by $|\mathbf{n},+n_{\mathrm{odd}}\rangle$). In contrast, the action $H_{-}H_{-}H_{+}$ on $\langle\mathbf{n}^{\prime}|$ increases the number of excitations by $n_{\mathrm{even}}=0,\pm 2,\pm 4\cdots$, or annihilated it (bases after action denoted by $\langle\mathbf{n}^{\prime},+n_{\mathrm{even}}|$). The reason is that any number of $H_{+}$ and an even number of $H_{-}$ do not change the parity of the total excitation numbers in $|\mathbf{n}\rangle$ or $\langle\mathbf{n}^{\prime}|$. Then any term after the actions of $H_{-}H_{+}$ and $H_{-}H_{-}H_{+}$ satisfies $\langle\mathbf{n}^{\prime\prime}|\hat{\rho}_{S}\otimes|\mathbf{n},+n_{\mathrm{odd}}\rangle\langle\mathbf{n}^{\prime},+n_{\mathrm{even}}|\mathbf{n}^{\prime\prime}\rangle=0$ where $\hat{\rho}_{S}$ is transformed from $\rho_{S}(t_{a})$. Similarly, any term as Eq. (54) with an odd number of $H_{-}$ turns $|\mathbf{n}\rangle$ and $\langle\mathbf{n}^{\prime}|$ into bases with different excitation numbers (due to different parities) or annihilate one or two of them. Then, any term obtained by expanding Eq. (54), if not annihilated, satisfies
$\langle\mathbf{n}^{\prime\prime}|\hat{\rho}_{S}\otimes|\mathbf{n},+n_{\mathrm{odd}}\rangle\langle\mathbf{n}^{\prime},+n_{\mathrm{even}}|\mathbf{n}^{\prime\prime}\rangle$ $=$ $0,$ (55)
$\mathrm{or}\ \langle\mathbf{n}^{\prime\prime}|\hat{\rho}_{S}\otimes|\mathbf{n},+n_{\mathrm{even}}\rangle\langle\mathbf{n}^{\prime},+n_{\mathrm{odd}}|\mathbf{n}^{\prime\prime}\rangle$ $=$ $0.$ (56)
So $\rho_{S}^{(l)}=0$ is guaranteed. On the other hand, if the number of $H_{-}$ is even in Eq. (54), there is
$K_{P}H_{\alpha}\!\cdots\!H_{\alpha}\rho_{S}(t_{a})\otimes|\mathbf{n}\rangle\langle\mathbf{n}^{\prime}|H_{\alpha}\!\cdots\!H_{\alpha}K_{P}$ (57)
$=$ $(-1)^{n_{\mathrm{even}}}H_{\alpha}\!\cdots\!H_{\alpha}K_{P}\rho_{S}(t_{a})\otimes|\mathbf{n}\rangle\langle\mathbf{n}^{\prime}|K_{P}H_{\alpha}\!\cdots\!H_{\alpha}$
$=$ $H_{\alpha}\!\cdots\!H_{\alpha}K_{P}\rho_{S}(t_{a})K_{P}\otimes|\mathbf{n}\rangle\langle\mathbf{n}^{\prime}|H_{\alpha}\!\cdots\!H_{\alpha}$
according to Eq. (51) and (52). Now we know that every term in Eq. (53) with an even number of $H_{-}$ satisfies Eq. (57) for any system initial state. By inserting Eq. (57) back into Eq. (53), $\mathcal{K}_{P}\bar{T}(t_{b},t_{a})=\bar{T}(t_{b},t_{a})\mathcal{K}_{P}$ is obtained.
The environment Hamiltonian with property (2) can be constructed by $H_{E}=\sum_{n\geqslant 1}H_{E_{n}}$ where each term $H_{E_{n}}$ consists of the product of a even number of ladder operators. For example, one or a collection of non-interacting (or interacting) harmonica oscillators. Similarly, the interaction Hamiltonian with property (3) can be constructed by $H_{S\!E}=\sum_{n\geqslant 1}S_{n}\otimes E_{n}$ where each $E_{n}$ consists of the product of an odd number of ladder operators. For example, the quantum Rabi interaction $\sigma^{+}a^{{\dagger}}+\sigma^{-}a+\sigma^{+}a+\sigma^{-}a^{{\dagger}}$ where the harmonic oscillator represents the environment.
V.2 Model and control
@@
In this subsection, we consider a quantum Rabi model where a qubit interacts with a dissipative harmonic oscillator in a Markovian reservoir at temperature $T$. The parity kicks are implemented on the qubit, i.e., the system (denoted by $S$), to protect it from the influence of the environment (the dissipative oscillator, denoted by $E$). The Hamiltonian of the qubit and oscillator is given by
@@
The parity kicks are chosen as $K_{P}=\sigma_{z}$, which is assumed to be ideal. It leads to an instantaneous phase change in the qubit’s quantum state. Starting from $t=0$, the controlled dynamics can be described by
@@
where $H_{1}(t)=\sum_{j=1}^{n}\delta(t-t_{j})H_{P}$ that realizes $K_{P}=\sigma_{z}$ at $t_{j}=j\tau$. It is easy to seen that conditions Eq. (5) and (6) are satisfied for our scheme. Therefore, the qubit can be dynamically decoupled from its environment after one or more cycles of parity kicks if $t_{c}$ is short enough. Starting from the condition $\rho_{S\!E}(0)=\rho_{S}(0)\otimes\rho_{E}^{\mathrm{th}}$ with arbitrary $\rho_{S}(0)$, the Choi-Jamiółkowski matrix of the target final dynamical map $T_{\mathrm{tar}}(t_{f},0)$ is given by
@@
In the following, we discuss the commutation relation $\mathcal{K}_{P}\bar{T}(t_{b},t_{a})=\bar{T}(t_{b},t_{a})\mathcal{K}_{P}$ for this model. The evolution of the qubit in the absence of control could be understood as $\rho_{S}(t_{b})=\bar{T}(t_{b},t_{a})\rho_{S}(t_{a})=\mathrm{Tr}_{E,R}[e^{-iH_{S\!E\!R}(t_{b}-t_{a})}\rho_{S}(t_{a})\otimes\rho_{E}^{\mathrm{th}}\otimes\rho_{R}^{\mathrm{th}}e^{iH_{S\!E\!R}(t_{b}-t_{a})}]$ where
@@
represents the Hamiltonian of the qubit, the oscillator, a Markovian reservoir (denoted by R) and their interactions. The oscillator and reservoir constitute the full environment of the qubit. The reservoir is in a thermal state at temperature $T$, which is given by
$\rho_{R}^{\mathrm{th}}=\prod_{j=1}^{\infty}\sum_{n_{j}=0}^{\infty}\frac{\bar{n}(\omega_{j})}{[1+\bar{n}(\omega_{j})]^{n_{j}+1}}|n_{j}\rangle\langle n_{j}|.$ (66)
Then the ERE\!-\!R initial state can be concisely expressed as
$\rho_{E}^{\mathrm{th}}\otimes\rho_{R}^{\mathrm{th}}$ $=$ $\sum\rho_{\mathbf{n}\mathbf{n}}|\mathbf{n}\rangle\langle\mathbf{n}|$ (67)
where $|\mathbf{n}\rangle=|n_{0},n_{1},n_{2}\cdots\rangle$, $p_{\mathbf{n}\mathbf{n}}$ is the diagonal element in the basis $|\mathbf{n}\rangle$. In this case, for every nonzero matrix element in Eq. (67), $n_{\mathrm{ket}}$ and $n_{\mathrm{bra}}$ always have the same parity since $n_{\mathrm{ket}}=n_{\mathrm{bra}}$. Now we divide the SERS\!-\!E\!-\!R Hamiltonian Eq. (65) into $H_{+}=H_{S}+H_{E}+H_{R}+H_{E\!R}$ and $H_{-}=H_{S\!E}$ such that $\sigma_{z}H_{+}=H_{+}\sigma_{z}$ and $\sigma_{z}H_{-}=-H_{-}\sigma_{z}$. Note that $H_{+}$ conserves the total excitation numbers in $|\mathbf{n}\rangle$ ($\langle\mathbf{n}|$) and thus preserves the parities of $n_{\mathrm{ket}}$ and $n_{\mathrm{bra}}$. In contrast, $H_{-}$ increases or decreases the total excitation number in $|\mathbf{n}\rangle$ ($\langle\mathbf{n}|$) by 1 and thus changes the parities of $n_{\mathrm{ket}}$ and $n_{\mathrm{bra}}$. Therefore, $\mathcal{K}_{P}\bar{T}(t_{b},t_{a})=\bar{T}(t_{b},t_{a})\mathcal{K}_{P}$ holds for this model according to the analysis in Sec. V.1. Besides, the commutation relation can be straightforwardly verified by numerically simulating Eq. (59) before and after the kick (with different qubit initial states), or directly calculating and comparing the Choi-Jamiółkowski matrices $\rho_{\mathcal{K}_{P}\bar{T}(t_{b},t_{a})}$ and $\rho_{\bar{T}(t_{b},t_{a})\mathcal{K}_{P}}$ by simulations.
@@
Following similar analysis or through numerical simulations, it is able to verify that the commutation relation might also holds for more related models. For example, a dissipative (or undamped) JC model, our model where the oscillator’s initial state satisfies $\rho_{E}(t_{I})=\sum_{n_{0}=0}^{\infty}p_{n_{0}}|n_{0}\rangle\langle n_{0}|$ (not limited to thermal states), a qubit coupled to multiple dissipative (or undamped) oscillators, a qubit coupled to a set of interacting oscillators, etc. Moreover, as known from the pseudomode theory [47, 48] and its recent progresses [51, 50, 49], the reduced dynamics of an open quantum system with a complex environment (infinite modes) is equivalent to that with a simpler environment (finite modes each subject to a Lindblad evolution) under certain conditions. Therefore, it is possible to apply our theory to more complex environments [52] to study the role of memory effects in parity kicks.